# Existence of cohomogeneity one Einstein metrics

### v4

@@ 3. Local Analysis
$p_{0}^{\pm}=\left(\pm\frac{1}{d_{1}},0,\frac{1}{d_{1}},0\right).$
@@
$\begin{bmatrix}\frac{4d_{2}-2d_{1}(d_{2}-1)}{d_{1}n}&\frac{d_{2}(d_{1}-1)(d_{1}-d_{2})}{d_{1}^{2}n}&\frac{2(d_{1}-1)}{d_{1}}&0\\ \frac{2d_{1}}{n}&\frac{1}{d_{1}}+\frac{d_{2}-d_{1}}{n}&0&\frac{d_{2}-1}{d_{1}}\\ \frac{2d_{2}}{d_{1}n}&\frac{d_{2}(d_{2}-d_{1})}{d_{1}^{2}n}&0&0\\ 0&0&0&\frac{2}{d_{1}}\end{bmatrix}$
The critical point is hyperbolic. The following are the only two unstable eigenvectors that are tangent to $\mathcal{E}$. Both eigenvectors have the same eigenvalue as $\frac{2}{d_{1}}$.
$v_{1}=\begin{bmatrix}-2d_{2}(d_{2}-1)\\ 2d_{1}(d_{2}-1)\\ -d_{2}(d_{2}-1)\\ 2d_{1}(d_{1}+1)\end{bmatrix},\quad v_{2}=\begin{bmatrix}-d_{1}^{2}+d_{1}d_{2}-n\\ -2d_{1}^{2}\\ -d_{2}\\ 0\end{bmatrix}$
From the unstable version of [CL55, Theorem 4.5, Chapter 13], there exists a continuous $1$-parameter family $\gamma_{s}$ that emanates $p_{0}^{+}$ such that
(3.1) $\gamma_{s}\sim p_{0}^{+}+se^{\frac{2}{d_{1}}\eta}v_{1}+e^{\frac{2}{d_{1}}\eta}v_{2}+O\left(e^{\left(\frac{2}{d_{1}}+\delta\right)\eta}\right).$
The integral curve $\gamma_{0}$ is in the invariant set $\mathcal{E}\cap\{Z=0\}$. Each $\gamma_{s}$ with $s>0$ is in the interior :=$\mathcal{E}^{\circ}:=\mathcal{E}\cap\{Y,Z>0\}\cap\{H^{2}<1\}$ and it represents a locally defined positive Einstein metric around $\mathcal{Q}^{d_{2}}$. As $v_{1}$ is tangent to $\mathcal{RF}^{+}$, the integral curve that emanates from $p_{0}^{+}$ along $v_{1}$ represents a locally defined Ricci-flat metric around $\mathcal{Q}^{d_{2}}$.
To show the existence of a complete Einstein metric on $M$ is equivalent to showing a $\gamma_{s}$ joins $p_{0}^{\pm}$ for some $s>0$. The existence of such a heterocline depends on the existence of critical points with non-zero $Y$ and $Z$, which satisfy the quadratic equation
@@
there are two real roots $\mu_{1}>\mu_{2}$ of the quadratic equation and hence two $\mathsf{G}$-invariant Einstein metrics on $\mathsf{G}/\mathsf{K}$. Subsequently, the following critical points exist.
$q_{i}^{\pm}=\left(\pm\frac{1}{n},\pm\frac{1}{n},y_{i},z_{i}\right),\quad R_{1}(y_{i},z_{i})=R_{2}(y_{i},z_{i})=\frac{n-1}{n^{2}},\quad i=1,2.$
Note that $\mu_{i}=\frac{z_{i}}{y_{i}}$. As $\mu_{1}$ and $\mu_{2}$ are roots for (3.2), we immediately obtain the following identities:
@@
Correspondingly, there exist two sine cone solutions of the cohomogeneity one Einstein equation that joins $q_{i}^{\pm}$, which are identified as the invariant sets
@@
(3.5) $n\leq 8,\quad A<\frac{(9-n)(d_{2}n+7n+9d_{1})}{(d_{1}n-8n-9d_{1})^{2}}\frac{d_{2}(d_{2}-1)^{2}}{4(d_{1}-1)}.$
For a $\mathsf{G}/\mathsf{K}$ that satisfies (3.5), one exploits the $\mathbb{Z}_{2}$-symmetry of the dynamical system to prove the existence of infinitely many cohomogeneity one Einstein metrics on $M$; see [Böh98, Corollary 5.8]. Unfortunately, there is no known example of a $\mathsf{G}/\mathsf{K}$ that satisfies (3.5) with $A>0$. We are hence motivated to find a new upper bound for $A$, presumably in the interval $\left(\frac{(9-n)(d_{2}n+7n+9d_{1})}{(d_{1}n-8n-9d_{1})^{2}}\frac{d_{2}(d_{2}-1)^{2}}{4(d_{1}-1)},\frac{1}{n+d_{1}}\frac{d_{2}(d_{2}-1)^{2}}{4(d_{1}-1)}\right)$, for the existence theorem.
@@
If $A\geq\frac{1}{n+d_{1}}\frac{d_{2}(d_{2}-1)^{2}}{4(d_{1}-1)}$, there exists at most one $\mathsf{G}$-invariant Einstein metric on $\mathsf{G}/\mathsf{K}$. It is further deduced that the two summands type ansatz does not yield any cohomogeneity one Einstein metric on $M$ if such an inequality holds; see [Böh98, Theorem 3.1]. In an upcoming paper [Chi25a], we show that for some principal orbit, its characteristic triple satisfies (3.3) yet does not yield any Einstein metric on $M$ from the two summands type ansatz.
@@ 4. Global Analysis
In this section, we assume (3.3) so that the two real roots $\mu_{1}>\mu_{2}$ exist. Our global analysis looks for a new upper bound for $A$ to guarantee the existence of a $\gamma_{s}$ that joins $p_{0}^{\pm}$. From [Böh98, Theorem 4.6], the existence of such a heterocline is implied if a $\gamma_{s}$ has a large enough winding angle around $\Phi_{2}$.
@@
By (3.4), we have $A=\frac{d_{2}(d_{2}-1)}{(n+d_{1})(\mu_{1}+\mu_{2})}\leq\frac{d_{2}(d_{2}-1)}{d_{1}\mu_{1}}$. The coefficient of $Z^{2}$ is positive. Since $H^{2}\leq 1$, each variable is bounded in $\mathcal{E}^{+}$ by (2.10). ∎
By (2.12), it is clear that an integral curve in $\mathcal{E}^{+}$ must escape the compact set through one of the three faces from the defining inequalities. Furthermore, by the derivatives of these defining functions, if an integral curve is not in $\mathcal{E}\cap\{Z=0\}$ or $\Phi_{i}$, it intersects $\partial\mathcal{E}^{+}$ transversely. Obviously, each $\gamma_{s}$ is initially in $\mathcal{E}^{+}$. Let $\sharp C(\gamma_{s})$ be the number of times that a $\gamma_{s}$ intersects
+$\mathcal{E}^{+}\cap\{H>0\}\cap\{X_{1}-X_{2}=0\}.$
We have the following proposition, which is an analogous statement to [Böh98, Theorem 4.6].
@@
on $\gamma_{0}$. Hence, the function $X_{1}-X_{2}$ is positive along $\gamma_{0}$, and we have $C(\gamma_{0})=0$.
Each $\gamma_{s}$ with $s\in(0,s_{\bullet})$ escapes $\mathcal{E}^{+}$ transversely. Let $\gamma_{s}(\eta_{s})$ denote the escaping point, we have the following connected path
@@
that joins a point in +$\mathcal{E}^{+}\cap\{H=0\}\cap\{X_{1}-X_{2}>0\}$ (from $\gamma_{0}$) and a point in +$\mathcal{E}^{+}\cap\{H>0\}\cap\{X_{1}-X_{2}=0\}$ (from $\gamma_{s_{\bullet}}$). If $\mathcal{P}$ does not intersect +$\mathcal{E}^{+}\cap\{X_{1}=X_{2}=0\}$, the set necessarily intersects the integral curve $\Phi_{1}$, which is a contradiction. Therefore, there exists a $s_{\star}\in(0,s_{\bullet})$ for which $\gamma_{s_{\star}}$ escapes $\mathcal{E^{+}}$ through +$\mathcal{E}^{+}\cap\{X_{1}=X_{2}=0\}$. By the $\mathbb{Z}_{2}$-symmetry of (2.9), such a $\gamma_{s_{\star}}$ is a heterocline that joins $p_{0}^{\pm}$. ∎
In [Chi24], we prove the existence of $s_{\bullet}$ by constructing a compact barrier set $\mathcal{S}$ that forces each $\gamma_{s}$ in the set to intersect +$\mathcal{E}^{+}\cap\{H>0\}\cap\{X_{1}-X_{2}=0\}$. With a different structural triple $(d_{1},d_{2},A)$, the set $\mathcal{S}$ may or may not exist. The counterpart to $\mathcal{S}$ in this paper depends on a varying $(d_{1},d_{2},A)$. As $A$ falls below or exceeds certain thresholds determined by $(d_{1},d_{2})$, each $\gamma_{s}$ in the set is forced to escape through certain faces.
@@
(4.3) $\begin{split}S^{\prime}&=-\frac{1}{X_{2}\left(X_{1}+\frac{d_{2}}{2d_{1}}X_{2}\right)}SP\\ T^{\prime}&=2TX_{2}\\ P^{\prime}&=P\left(H\left(3G+\frac{3}{n}(1-H^{2})-1\right)+\frac{n}{d_{1}}X_{2}-X_{1}\right)+(X_{1}-X_{2})(P+Q),\end{split}$
@@
By Proposition 4.1, the set $\mathfrak{S}_{A}(\sigma,\tau)$ is compact.
@@
Suppose a non-trivial integral curve escapes through the face $\{X_{2}=0\}$. If $X_{1}=0$ at the escaping point, there is nothing to prove. If $X_{1}>0$ at the escaping point, the $Z$-coordinate must vanish as the function $S$ is non-increasing in $\mathfrak{S}_{A}(k,l)$. With $X_{2}=Z=0$, the inequality $P\geq 0$ becomes $-X_{1}\frac{1-d_{1}^{2}X_{1}^{2}}{n}\geq 0$, meaning that the escaping point can only be $p_{0}^{+}$, a contradiction.
@@
The inequality $S\leq\sigma<\frac{n+d_{1}}{2d_{1}}\mu_{1}$ in $\mathfrak{S}_{A}(\sigma,\tau)$ implies
$\frac{Z}{Y}\frac{n+d_{1}}{2d_{1}}<\frac{Z}{Y}\frac{X_{1}+\frac{d_{2}}{2d_{1}}X_{2}}{X_{2}}=S<\frac{n+d_{1}}{2d_{1}}\mu_{1}.$
Therefore, an integral curve in $\mathfrak{S}_{A}(\sigma,\tau)$ does not escape through the face $\{Z-\mu_{1}Y=0\}$.

### v5

@@ 3. Local Analysis
(3.1) $p_{0}^{\pm}=\left(\pm\frac{1}{d_{1}},0,\frac{1}{d_{1}},0\right).$
@@
$\begin{bmatrix}\frac{4d_{2}-2d_{1}(d_{2}-1)}{d_{1}n}&\frac{d_{2}(d_{1}-1)(d_{1}-d_{2})}{d_{1}^{2}n}&\frac{2(d_{1}-1)}{d_{1}}&0\\ \frac{2d_{1}}{n}&\frac{1}{d_{1}}+\frac{d_{2}-d_{1}}{n}&0&\frac{d_{2}-1}{d_{1}}\\ \frac{2d_{2}}{d_{1}n}&\frac{d_{2}(d_{2}-d_{1})}{d_{1}^{2}n}&0&0\\ 0&0&0&\frac{2}{d_{1}}\end{bmatrix}$.
The critical point is hyperbolic. The following are the only two unstable eigenvectors that are tangent to $\mathcal{E}$. Both eigenvectors have the same eigenvalue $\frac{2}{d_{1}}$,
$v_{1}=\begin{bmatrix}-2d_{2}(d_{2}-1)\\ 2d_{1}(d_{2}-1)\\ -d_{2}(d_{2}-1)\\ 2d_{1}(d_{1}+1)\end{bmatrix},\quad v_{2}=\begin{bmatrix}-d_{1}^{2}+d_{1}d_{2}-n\\ -2d_{1}^{2}\\ -d_{2}\\ 0\end{bmatrix}$.
By the Hartman–Grobman theorem, there is a 1-to-1 correspondence between each choice of linearized solution and an actual solution curve that emanates from $p_{0}^{+}$. From the unstable version of [CL55, Theorem 4.5, Chapter 13], the error of such a correspondence is $O\left(e^{\left(\frac{2}{d_{1}}+\delta\right)\eta}\right)$. There exists a continuous $1$-parameter family $\gamma_{s}$ that emanates from $p_{0}^{+}$ such that
(3.2) $\gamma_{s}\sim p_{0}^{+}+(sv_{1}+v_{2})e^{\frac{2}{d_{1}}\eta}+O\left(e^{\left(\frac{2}{d_{1}}+\delta\right)\eta}\right).$
As $v_{1}$ and $v_{2}$ are tangent to $\mathcal{E}$, each $\gamma_{s}$ is in $\mathcal{E}$.
By (2.13) and the linearized solution, we have
$\underline{f}^{2}=\lim\limits_{t\to 0}f_{2}^{2}=\lim\limits_{\eta\to-\infty}\frac{1-H^{2}}{n\Lambda YZ}=\frac{d_{1}}{\Lambda s}$
along the integral curve $\gamma_{s}$. Eigenvectors $v_{1}$ and $v_{2}$ separately correspond to two extreme situations. As $v_{1}$ is tangent to $\mathcal{RF}^{+}$, the integral curve that emanates from $p_{0}^{+}$ along $v_{1}$ represents a locally defined Ricci-flat metric around $\mathcal{Q}^{d_{2}}$. The vector $v_{2}$ tangent to $\{Z=0\}$ and it leads to an integral curve lying in $\{Z=0\}$, where the metric component along $\mathcal{Q}^{d_{2}}$ diverges. We restrict to directions of the form $sv_{1}+v_{2}$ rather than the more general combination $s_{1}v_{1}+s_{2}v_{2}$ since scalar multiplication does not yield genuinely new metrics: it simply shifts the integral curve along the $\eta$-axis. In the original coordinate, the operation corresponds to a homothetic change of Ricci-flat metrics, or translation in the $t$-axis of positive Einstein metrics.
By (3.2), we have
(3.3) $Z(\gamma_{s})\sim 2d_{1}(d_{1}+1)e^{\frac{2}{d_{1}}\eta}+O\left(e^{\left(\frac{2}{d_{1}}+\delta\right)\eta}\right),\quad H(\gamma_{s})\sim 1-d_{1}n(d_{1}+1)e^{\frac{2}{d_{1}}\eta}+O\left(e^{\left(\frac{2}{d_{1}}+\delta\right)\eta}\right)$
near $p_{0}^{+}$. Therefore, the essential 1-parameter family of non-degenerate locally defined positive Einstein metrics around $\mathcal{Q}^{d_{2}}$ is represented by $\{\gamma_{s}\mid s>0\}$. Since the $Y$-coordinate of $p_{0}^{+}$ is positive, each $\gamma_{s}$ with $s>0$ is in the interior
$\mathcal{E}^{\circ}:=\mathcal{E}\cap\{Y,Z>0\}\cap\{H^{2}<1\}.$
With (3.2), it is also natural to let $\gamma_{\infty}$ denote the integral curve that emanates from $p_{0}^{+}$ along $v_{1}$ and stays in $\mathcal{RF}^{+}$.
To obtain a complete Einstein metric on $M$, one needs to show the existence of an integral curve that joins $p_{0}^{\pm}$. In Böhm’s analysis for $A=0$, such heteroclines arise because integral curves emanating from $p_{0}^{+}$ spiral around the sine cone based over a homogeneous Einstein metric on $\mathsf{G}/\mathsf{K}$. In our setting with $A>0$, a similar mechanism requires the existence of homogeneous Einstein metrics on $\mathsf{G}/\mathsf{K}$. Algebraically, this corresponds to critical points that are characterized by the condition $X_{1}=X_{2}=\frac{1}{n}$ and $R_{1}=R_{2}$, which reduces to the following quadratic equation,
@@
there are two real roots $\mu_{1}>\mu_{2}$ of the quadratic equation. These roots correspond to metric ratios $\frac{f_{1}^{2}}{f_{2}^{2}}$ that yield two distinct homogeneous Einstein metrics on $\mathsf{G}/\mathsf{K}$, represented by the following critical points in $\mathcal{RF}^{\pm}$,
$q_{i}^{\pm}=\left(\pm\frac{1}{n},\pm\frac{1}{n},y_{i},z_{i}\right),\quad\frac{z_{i}}{y_{i}}=\mu_{i},\quad R_{1}(y_{i},z_{i})=R_{2}(y_{i},z_{i})=\frac{n-1}{n^{2}},\quad i=1,2.$
Being treated as trivial integral curves on $\mathcal{RF}^{+}$, the points $q_{i}^{+}$ also represent the Ricci-flat cones over the two homogeneous Einstein metrics, while $q_{i}^{-}$ are their time reversal counterparts related by the $\mathbb{Z}_{2}$-symmetry.
As $\mu_{1}$ and $\mu_{2}$ are roots for (3.4), we immediately obtain the following identities:
(3.6) $R_{1}-R_{2}=\frac{n+d_{1}}{d_{2}}A(Z-\mu_{1}Y)(Z-\mu_{2}Y),$
where
@@
The two sine cone solutions of the cohomogeneity one Einstein equation that joins $q_{i}^{\pm}$ are identified as the invariant sets
@@
(3.8) $n\leq 8,\quad A<\frac{(9-n)(d_{2}n+7n+9d_{1})}{(d_{1}n-8n-9d_{1})^{2}}\frac{d_{2}(d_{2}-1)^{2}}{4(d_{1}-1)},$
so there exist integral curves emanating from $p_{0}^{+}$ that spiral around the cone with arbitrarily large winding number, which by the $\mathbb{Z}_{2}$-symmetry leads to infinitely many heteroclines that join $p_{0}^{\pm}$. Hence there exist infinitely many cohomogeneity one Einstein metrics on $M$; see [Böh98, Corollary 5.8]. Unfortunately, there is no known example of a $\mathsf{G}/\mathsf{K}$ that satisfies (3.8) with $A>0$. We are hence motivated to find a new upper bound for $A$, presumably in the interval $\left(\frac{(9-n)(d_{2}n+7n+9d_{1})}{(d_{1}n-8n-9d_{1})^{2}}\frac{d_{2}(d_{2}-1)^{2}}{4(d_{1}-1)},\frac{1}{n+d_{1}}\frac{d_{2}(d_{2}-1)^{2}}{4(d_{1}-1)}\right)$, for the existence theorem.
@@
If $A\geq\frac{1}{n+d_{1}}\frac{d_{2}(d_{2}-1)^{2}}{4(d_{1}-1)}$, there exists at most one $\mathsf{G}$-invariant Einstein metric on $\mathsf{G}/\mathsf{K}$. It is further deduced that the two summands type ansatz does not yield any cohomogeneity one Einstein metric on $M$ if such an inequality holds; see [Böh98, Theorem 3.1]. In an upcoming paper [Chi25b], we show that for some principal orbit, its structural triple satisfies (3.5) yet does not yield any Einstein metric on $M$ from the two summands type ansatz.
@@ 4. Global Analysis
As explained in the Introduction, Böhm’s local spiraling mechanism (which yields infinitely many solutions when $q_{2}^{\pm}$ are foci) is generally not available if $A>0$. In this section, we implement a global analysis outlined below: construct suitable barrier functions, single out a distinguished integral curve $\gamma_{s_{\bullet}}$, and use an intermediate value argument to produce a heterocline joining $p_{0}^{\pm}$. We begin by introducing the compact region $\mathcal{E}^{+}$ and establishing its basic properties. We then define the barrier functions $S,T,P$ and analyze their evolution along $\gamma_{s}$.
@@
By (3.7), we have $A=\frac{d_{2}(d_{2}-1)}{(n+d_{1})(\mu_{1}+\mu_{2})}\leq\frac{d_{2}(d_{2}-1)}{d_{1}\mu_{1}}$. The coefficient of $Z^{2}$ in (4.1) is positive. Therefore, the functions $X_{1}-X_{2}$, $Y$ and $Z$ are bounded by (2.10). As $H^{2}\leq 1$ in $\mathcal{E}$, the functions $X_{1}$ and $X_{2}$ are bounded in a compact parallelogram region. The proof is complete. ∎
Proposition 4.2.
Every integral curve in the interior
$(\mathcal{E}^{+})^{\circ}:=\mathcal{E}^{\circ}\cap\{H>0\}\cap\{X_{1}-X_{2}>0\}\cap\{\mu_{1}Y-Z>0\}$
leaves $\mathcal{E}^{+}$ transversely through one of the following boundary faces
$\{H=0\},\quad\{X_{1}-X_{2}=0\},\quad\{\mu_{1}Y-Z\}.$
Proof.
By (2.12), an integral curve in $(\mathcal{E}^{+})^{\circ}$ must escape $\mathcal{E}^{+}$. Since $\mathcal{RF}^{+}$, $\mathcal{E}\cap\{Y=0\}$, and $\mathcal{E}\cap\{Z=0\}$ are invariant, an integral curve in $(\mathcal{E}^{+})^{\circ}$ can only exit $\mathcal{E}^{+}$ through one of the boundary faces listed above.
If an integral curve exits through $\{H=0\}$, transversality follows from (2.12). For $\{X_{1}-X_{2}=0\}$, differentiating gives
(4.2) $\begin{split}(X_{1}-X_{2})^{\prime}&=(X_{1}-X_{2})H\left(G+\frac{1-H^{2}}{n}-1\right)+R_{1}-R_{2}\\ &=(X_{1}-X_{2})H\left(G+\frac{1-H^{2}}{n}-1\right)+\frac{n+d_{1}}{d_{2}}A(Z-\mu_{1}Y)(Z-\mu_{2}Y),\end{split}$
A non-transverse exit would force the integral curve to lie entirely in the invariant sets $\mathcal{E}\cap\{Y=Z=0\}$ or in $\Phi_{i}$, both disjoint from $(\mathcal{E}^{+})^{\circ}$.
Similarly, for $\{\mu_{i}Y-Z=0\}$ one finds
(4.3) $(\mu_{i}Y-Z)^{\prime}=(\mu_{i}Y-Z)H\left(G+\frac{1-H^{2}}{n}\right)-\mu_{i}YX_{1}-Z(X_{1}-2X_{2}).$
So non-transverse escape is possible only in $\mathcal{E}\cap\{Y=Z=0\}$, or in $\Phi_{i}$. Again, these sets are disjoint from $(\mathcal{E}^{+})^{\circ}$.
Thus any integral curve in $(\mathcal{E}^{+})^{\circ}$ escapes $\mathcal{E}^{+}$ only transversely. ∎
By (3.3) and the coordinate of $p_{0}^{+}$, each $\gamma_{s}$ with $s>0$ is initially in $(\mathcal{E}^{+})^{\circ}$. Therefore, each $\gamma_{s}$ with $s>0$ escapes $\mathcal{E}^{+}$ transversely by Proposition 4.2.
Let $\sharp C(\gamma_{s})$ be the number of times that a $\gamma_{s}$ intersects
$\mathcal{B}:=\mathcal{E}^{+}\cap\{X_{1}-X_{2}=0\}\cap\{0<H<1\}.$
We have the following proposition.
Proposition 4.3.
If $\sharp C(\gamma_{s})\geq 1$ for some $s>0$, the integral curve escapes $\mathcal{E}^{+}$ through the 2-dimensional region
$\tilde{\mathcal{B}}:=\mathcal{E}\cap\{X_{1}-X_{2}=0\}\cap\{0<H<1\}\cap\{\mu_{2}Y<Z<\mu_{1}Y\}.$
Proof.
We show that the escaping point of $\gamma_{s}$ necessarily contributes to $\sharp C(\gamma_{s})$. By Proposition 4.2, we need to rule out the other two possibilities.
If a $\gamma_{s}$ escapes through $\{H=0\}$, the function $H$ becomes and remains negative after the intersection. Thus $\sharp C(\gamma_{s})=0$.
If a $\gamma_{s}$ escapes through $\{\mu_{1}Y-Z\}$, the integral curve enters the set
$\mathcal{U}:=\mathcal{E}\cap\{\mu_{1}Y-Z<0\}\cap\{X_{1}-X_{2}>0\}\cap\{Z>0\}.$
By (4.3), we have
$\left.(\mu_{1}Y-Z)^{\prime}\right|_{\mu_{1}Y-Z=0}=2Z(X_{2}-X_{1})<0.$
By (4.2) and $Z-\mu_{2}Y>Z-\mu_{1}Y>0$, we also have
$\left.(X_{1}-X_{2})^{\prime}\right|_{X_{1}-X_{2}=0}=\frac{n+d_{1}}{d_{2}}A(Z-\mu_{1}Y)(Z-\mu_{2}Y)>0.$
The set $\mathcal{U}$ is invariant. So it is impossible for a $\gamma_{s}$ in $\mathcal{U}$ to intersect $\mathcal{B}$ afterward.
The only possibility is $\gamma_{s}$ escaping $\mathcal{E}^{+}$ through $\{X_{1}-X_{2}=0\}$. As $X_{1}-X_{2}>0$ at $p_{0}^{+}$, we have
$\left.(X_{1}-X_{2})^{\prime}\right|_{X_{1}-X_{2}=0}=\frac{n+d_{1}}{d_{2}}A(Z-\mu_{1}Y)(Z-\mu_{2}Y)\leq 0$
at the intersection point. The derivative cannot vanish at the intersection point, since $\gamma_{s}$ does not intersect $\Phi_{i}$. Therefore, we conclude that such a $\gamma_{s}$ must intersect $\tilde{\mathcal{B}}$. ∎
We have the following lemma, which is an analogous statement to [Böh98, Theorem 4.6].
@@
on $\gamma_{0}$. Hence, the function $X_{1}-X_{2}$, which is initially positive, can never vanish along $\gamma_{0}$. We have $C(\gamma_{0})=0$.
Let $\eta_{s}\in\mathbb{R}$ be the smallest number such that $\gamma_{s}(\eta_{s})\in\partial\mathcal{E}^{+}$. The above argument shows that $\gamma_{0}(\eta_{0})\in\mathcal{E}^{+}\cap\{H=0\}\cap\{X_{1}-X_{2}>0\}$. By Proposition 4.3, we have $\gamma_{s_{\bullet}}(\eta_{s_{\bullet}})\in\tilde{\mathcal{B}}\subset\{X_{1}-X_{2}=0\}$. Since each $\gamma_{s}$ with $s>0$ escapes $\mathcal{E^{+}}$ transversely by Proposition 4.2, we have the following connected path
@@
that joins a point in $\mathcal{E}^{+}\cap\{H=0\}\cap\{X_{1}-X_{2}>0\}$ (from $\gamma_{0}$) and a point in $\tilde{\mathcal{B}}\subset\{X_{1}-X_{2}=0\}$ (from $\gamma_{s_{\bullet}}$). Since $\gamma_{0}(\eta_{0})$ is not in the closure of $\tilde{\mathcal{B}}$. The path $\mathcal{P}$ intersects $\partial\tilde{\mathcal{B}}$. Therefore, there exists a $s_{\star}\in(0,s_{\bullet})$ for which $\gamma_{s_{\star}}(\eta_{s_{\star}})\in\partial\tilde{\mathcal{B}}$. Since $\mathcal{RF}^{+}$ and each $\Phi_{i}$ are invariant, the point $\gamma_{s_{\star}}(\eta_{s_{\star}})$ must be in $\partial\tilde{B}\cap\{H=0\}\subset\{X_{1}=X_{2}=0\}$, meaning that $\gamma_{s_{\star}}$ escapes $\mathcal{E^{+}}$ through $\{X_{1}=X_{2}=0\}$. By the $\mathbb{Z}_{2}$-symmetry explained in Remark 2.2, such a $\gamma_{s_{\star}}$ is a heterocline that joins $p_{0}^{\pm}$. ∎
In [Chi24], we prove the existence of $s_{\bullet}$ by constructing a compact barrier set $\mathcal{S}$ that forces each $\gamma_{s}$ in the set to intersect $\mathcal{E}^{+}\cap\{H>0\}\cap\{X_{1}-X_{2}=0\}$. With a different structural triple $(d_{1},d_{2},A)$, the set $\mathcal{S}$ may or may not exist. The counterpart to $\mathcal{S}$ in this paper depends on varying $(d_{1},d_{2},A)$. As $A$ falls below or exceeds certain thresholds determined by $(d_{1},d_{2})$, each $\gamma_{s}$ in the set is forced to escape through certain faces.
@@
(4.5) $S^{\prime}$ $=-\frac{1}{X_{2}\left(X_{1}+\frac{d_{2}}{2d_{1}}X_{2}\right)}SP$
(4.6) $T^{\prime}$ $=2TX_{2}$
(4.7) $P^{\prime}$ $=P\left(H\left(3G+\frac{3}{n}(1-H^{2})-1\right)+\frac{n}{d_{1}}X_{2}-X_{1}\right)+(X_{1}-X_{2})(Q+P),$
@@
We discuss our motivation of using the functions listed in (4.4). In the original coordinates, we have $T=\Lambda f_{2}^{2}$. From the analytic viewpoint, both $T$ and its derivative (4.6) is positive in $(\mathcal{E}^{+})^{\circ}$, making $T$ a convenient function to control along the integral curves. The function $S$ is also positive in $(\mathcal{E}^{+})^{\circ}$, and its monotonicity depends on the sign of $P$. From the dynamical systems viewpoint, the function $S$ monitors the rotation of $\gamma_{s}$ around $\Phi_{2}$, which occurs in the $(X_{1}-X_{2},\mu_{2}Y-Z)$–plane, while $H$ decreases monotonically. The ratio $\frac{ZX_{1}}{YX_{2}}$ is a natural quantity to record this rotation, but its derivative involves a factor (the original version of $P$) that does not preserve positivity in $(\mathcal{E}^{+})^{\circ}$. The modified definition of $S$ is precisely tuned so that the derivative formulas (4.5)–(4.7) are better suited for our analysis.
From a geometric viewpoint, the function $S$ can be regarded as a perturbation of certain first integrals that arise in the Ricci–flat case. For instance, many known cohomogeneity one Ricci–flat metrics with a first integral (e.g. the Bryant–Salamon $G_{2}$ and $\mathrm{Spin}(7)$ metrics [BS89]) satisfy the following first order system in the original $(t,f_{1},f_{2})$-coordinates:
(4.9) $\frac{\dot{f_{1}}}{f_{1}}=\frac{1}{f_{1}}-\frac{d_{2}}{2d_{1}\sigma}\frac{f_{1}}{f_{2}^{2}},\quad\frac{\dot{f_{2}}}{f_{2}}=\frac{1}{\sigma}\frac{f_{1}}{f_{2}^{2}},\quad\sigma=\frac{d_{1}+1}{d_{2}-1}.$
In the new $(\eta,X_{1},X_{2},Y,Z)$-coordinates, this first integral is realized as an algebraic curve in $\mathcal{RF}^{+}$ given by
$\mathcal{L}:=\mathcal{RF}^{+}\cap\{X_{1}=Y-\tfrac{d_{2}}{2d_{1}\sigma}Z\}\cap\left\{X_{2}=\frac{Z}{\sigma}\right\}\cap\left\{\sigma=\frac{d_{1}+1}{d_{2}-1}\right\}.$
We refer to Appendix A for the proof of $\mathcal{L}$ being invariant for some $(d_{1},d_{2},A)$ and explicit examples where the integrability condition can be verified. Our definition of $S$ unifies the two algebraic relations in $\mathcal{L}$ into the single condition $S=\sigma$. Perturbing $(d_{1},d_{2},A,\sigma)$ away from the Ricci–flat subsystem, together with the functions $P$ and $T$, provides a useful barrier in the positive Einstein setting.
@@
The set (4.10) is a closed subset of the compact set $\mathcal{E}^{+}$, see Proposition 4.1, and it is therefore compact.
Remark 4.5.
In principle, other choices of $(\sigma,\tau)$ are possible to obtain an upper bound of existence for $A$. The present choice is artificially tuned so that $\gamma_{s_{\bullet}}$, where
(4.11) $s_{\bullet}=\frac{d_{1}}{\tau}=\frac{d_{1}\sigma}{(d_{2}-1)\sigma-(d_{1}+1)},$
is the only candidate in the 1-parameter family to enter the set $\mathfrak{S}_{A}(\sigma,\tau)$ that plays the role in Lemma 4.4. Furthermore, we show in [Chi25b] that the present choice of $(\sigma,\tau)$ maximizes the admissible upper bound for $A$ under the framework of this paper.
Here and in what follows, we take $s_{\bullet}$ as in (4.11). With this preparation, we first analyze the possible possible faces of $\mathfrak{S}_{A}(\sigma,\tau)$, through which an integral curve inside $\mathfrak{S}_{A}(\sigma,\tau)$ escapes. Next, we show that for suitable values of $A$, the integral curve $\gamma_{s_{\bullet}}$ indeed enters $\mathfrak{S}_{A}(\sigma,\tau)$ once it leaves $p_{0}^{+}$. Lastly, we identify ranges of $A$ for which escape through $\mathfrak{S}_{A}(\sigma,\tau)\cap\{P=0\}$ is impossible, forcing $\gamma_{s_{\bullet}}$ to leave through $\mathfrak{S}_{A}(\sigma,\tau)\cap\{X_{1}=X_{2}>0\}$ and hence producing the desired heterocline.
@@
Suppose a non-trivial integral curve escapes through the face $\{X_{2}=0\}$. If $X_{1}=0$ at the escaping point, there is nothing to prove since the escape point lies in $\{P=0\}$. If $X_{1}>0$ at the escaping point, the $Z$-coordinate must vanish as the function $S$ is non-increasing in $\mathfrak{S}_{A}(k,l)$. With $X_{2}=Z=0$, the inequality $P\geq 0$ becomes $X_{1}\frac{1-d_{1}^{2}X_{1}^{2}}{n}\geq 0$, meaning that the escaping point can only be $p_{0}^{+}$, a contradiction.
@@
By $X_{1}\geq X_{2}$ and $n=d_{1}+d_{2}$, we also have
$\frac{Z}{Y}\frac{n+d_{1}}{2d_{1}}<\frac{Z}{Y}\frac{X_{1}+\frac{d_{2}}{2d_{1}}X_{2}}{X_{2}}=S.$
Therefore, the inequality $S\leq\sigma<\frac{n+d_{1}}{2d_{1}}\mu_{1}$ in $\mathfrak{S}_{A}(\sigma,\tau)$ implies
$\frac{Z}{Y}\frac{n+d_{1}}{2d_{1}}<\frac{n+d_{1}}{2d_{1}}\mu_{1}.$
Hence, an integral curve in $\mathfrak{S}_{A}(\sigma,\tau)$ does not escape through the face $\{Z-\mu_{1}Y=0\}$.